how many grams of 5% antiseptic solution must be added to 300g of a 10% solution to produce an 8% solution?
Answer by stanbon(75887)

how many grams of 5% antiseptic solution must be added to 300g of a 10% solution to produce an 8% solution?
----
Eqution:
anti + anti = anti
0.05x + 0.10*300 = 0.08(x+300)
-----------------------------------
Multiply thru by 100 to get:
5x + 10*300 = 8x+8*300
3x = 2*300
x = 200 g (amt. of 5% solution needed)
====================
Cheers,
Stan H.
====================